1. Using the method described in the lab, you wanted to determine the formula of the transition metal oxide, zinc oxide. Starting with 3.145 g of zinc you formed 3.915 g of oxide. Determine the mass of oxygen in this compound. Express your answer in grams as three significant digits. 2. Is this compound (circle one): a. covalent b. elemental c. ionic d. undeterminable 3. Calculate the empirical formula and molecular formula of the compound. Calculate the percentage by mass of the elements in the compound. Express your answers as three significant digits. 4.
Answer #1

Q1.

mass of oxygen = (Mass of oxide ) - (Metal) = 3.915-3.145 = 0.770 g

Q2

the compound (oxide) must be ionic, since Zn+2 and O-2 are charged

Q3

empirical formula

mol of Zn = mass/MW = 3.145/65.39 = 0.048

mol of O = mass/MW = 0.77/16 = 0.048125

raito 0.048125 /0.048 = 1

Zn:O

ZnO

Q4.

% by mass of Zn = mass of Zn /Y Total mass * 100 = 3.145/3.915*100 = 80.33%

% mass of O = 100-80.33 = 19.67% for O